PRECISE DETERMINATION OF THE B AND C STATES OF XeF FROM A DEPERTURBATION ANAL-YSIS OF THE B → X SPECTRUM OF ^{136} XeF

K. M. LUCKETT, P. C. TELLINGHUISEN, and \underline{J. TELLINGHUISEN}, Department of Chemistry, Vanderbilt University, Nashville, TN 37235.